OpenStudy (anonymous):

Tickets to a show cost different for seniors, children, and adults. Adult tickets are twice as much as senior tickets. Two senior tickets plus one children’s ticket cost $17. Two children tickets plus one adult ticket cost $24. How much would it cost to buy one adult ticket and one senior ticket?

OpenStudy (bahrom7893):

Let's make a system of equations: s - senior, c - children, a - adults

OpenStudy (bahrom7893):

Adult tickets are twice as much as senior tickets. a = 2*s Two senior tickets plus one children’s ticket cost $17 2*s + c = 17 Two children tickets plus one adult ticket cost $24. 2*c + a = 24 So: a = 2s 2s + c = 17 2c + a = 24

OpenStudy (bahrom7893):

You're looking for a+s, so find a and s.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

still dont get the question

OpenStudy (anonymous):

it a 9th grade grade question it confusing